Charles has to pay $103 per year for the next 6 years. If the appropriate discount rate is 0.128, what would it cost today to settle his debt?
Answer #1

This question requires application of PV of annuity formula, according to which

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

PV = 103 * 1-(1+0.128)-6 PV = 103* 0.128

PV = 103 * 4.0199

PV = $414.05